OpenStudy (anonymous):

let a,b,c belongs to R. using the properties of real numbers if ab=1 prove that b=1/a

OpenStudy (agreene):

Given: \[a,b,c \in \mathbb{R}\] \[ab=1\] using the properties of addition: \[a^{-1}*ab=1*a^{-1}\] \[\cancel{a^{-1}}\cancel{a}b=1*a^{-1}\] \[b=a^{-1}\] which is the same as: \[b=\frac{1}{a}\] from the definition of exponentiation (which is a special form of multiplication--which is a special form of addition, which is why we only need to use the properties of addition) \[QED\]

OpenStudy (anonymous):

thanks for your help